graph the solution on the number line
-1/5|-2x+3|=-3
*(less than or equal to -3)

Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

-1/5 * |-2x + 3| = -3
|-2x + 3| = 3/5

Case 1: -2x + 3 = 3/5
-2x = -12/5
x = 6/5

Case 2: -2x + 3 = -3/5
-2x = -18/5
x = 9/5

1......6/5......7/5......8/5......9/5......10
<-----|------------------------|------>

Answer by ikleyn(52794)

.

Two sentences that are written in this post, are INCOMPATIBLE one with the other,

and, considered as one whole thing, MAKE no SENSE.


If they are two separate sentences, then separate them explicitly in the post
and DO NOT FORCE the reader to guess what they mean together.


The "solution" by @CubeyThePenguin is nonsense, too.